# Application of AI-based Models for Online Fraud Detection and Analysis

### v1

@@ 2.1.3 Data extraction
Next, the authors agreed on the data that should be extracted from the included studies. The task of extracting data was carried out by only one of the three reviewers. This was deemed sufficient since the reviewer’s role involved extracting the required details from the papers, and it was unnecessary for a second reviewer to check accuracy. The only aspect of the data extraction that the reviewer had to conceptualize was the specific Fraud Type analyzed by the study under question. For example, if a study is analyzing Phishing URLs, then it is labelled as Phishing URLs.
Not all papers explicitly specified the type of fraud they were analysing. Due to the diversity of scams, there is no agreed way of labelling fraud types. As as result, we used an umbrella term to categorise them. For example, online scam campaigns, made by bot users to advertise fraudulent phishing URLs include fake users and phishing URLs fraud analysis, hence we agreed to label papers of broad online scam campaigns as social media scams.
To ease the representation of our findings, we do not classify a paper with more than one fraud type. More specifically, the reviewer had to resort to the main goal of the paper towards annotating the kind of fraud it analyses. For example, if a paper used phishing emails to extract phishing URLs towards detecting phishing URLs, then that paper would be labelled as Phishing URLs, as that is the main goal of the study. The final version of the data extracted from each record is depicted in Table 2, along with relevant examples.
@@ 3.1 Study selection and characteristics
The PRISMA-ScR flow diagram in Figure 3 summarises the study selection process for the two academic searches. $2,617$ stuides were identified for eligibility screening. The ACM Digital Library returned $389$ documents, IEEE Xplore returned $712$ documents, Web of Science returned $253$ documents, ProQuest returned $783$ documents, Google Scholar returned $399$ documents, and ArXiv returned $47$ documents. An additional $15$ papers were recommended by the experts in the area. After removing duplicates, $2,457$ papers remained for further review.
@@
At this stage, $10\%$ of these papers were selected ($N=242$) for Inter-Rater Reliability, to calculate the multi-annotator agreement between the three annotators of this review. The Fleiss Kappa score between all three annotators was $0.83$, indicating almost perfect agreement. The Cohen Kappa Agreement was also calculated between each pair of annotators. The agreement between the annotator A and the annotator B was $0.65$ (substantial agreement), between A and annotator C was $0.66$, and between B and C was $0.52$ (moderate agreement). [^1] Annotators A, B, and C will be replaced with the initials of authors after the reviewing process. The three annotators compared their annotation process and reviewed the eligibility criteria and goals of this review. Following this discussion, the lead annotator moved to perform the rest of the annotation of the papers to be included in this review.
Reviewing the abstract of those papers resulted in $2,107$ papers being excluded from the study as they did not fit the eligibility criteria discussed in Section 2.1.2. This resulted in $350$ full-papers passing to eligibility screening, out of which $127$ did not fit the eligibility criteria and was excluded. Overall, this process resulted in $223$ full-text papers being included for qualitative analysis.
@@
3.2 Types of Online Fraud
@@
Figure 4 shows the types of fraud analyzed in the full-text papers that were included in the qualitative analysis. The reviewer of these studies manually coded each paper with the scam type that the study is focusing on, based on the title, abstract, and methodology of the studies. The majority of studies that were included for qualitative analysis focus on phishing detection, with about a third ($29\%$) of the studies focusing on detecting phishing URLs online ($N=64$). More specifically, these works tackle the problem of automatically detecting whether a given URL is likely to be fraudulent. A large number of papers were related to detecting phishing emails ($N=29$), followed by studies on SMS phishing detection ($N=20$). Other studies on phishing include phone call transcripts towards understanding and detecting voice call phishing ($N=12$) and a small number of studies attempt to understand phishing methods via victim reports ($N=4$).
@@
Moving on to other types of fraud, we found many studies that attempted to detect fake reviews on various platforms like Google Play Store, Apple App Store, Yelp, and TripAdvisor ($N=23$). Another widely studied scam focus was recruitment fraud ($N=20$). We also found a number of studies that employed AI techniques to detect fake accounts on Facebook, Instagram, and Twitter ($N=18$).
A few studies used Generative AI (GenAI) to gain a better understanding of some of the latest methods of phishing and other fraud types. GenAI represents cases where advanced LLMs are misused towards social engineering attacks. GenAI models, have revolutionised various industries as they are now able to promptly generate human-like text in response to input commands. $3$ studies employed GenAI models to automatically interact with scammers to waste the scammers’ resources or time, and to collect information regarding the methods with which the scammers try to defraud users to disrupt their operation. This is defined as scam baiting, the process of using generative AI models to deceive and engage with scammers. Interestingly, this is a countermeasure against online fraud (GenAI Scam Bait in the figure).
Our search also returned many studies that discuss the misuses of GenAI towards Social Engineering Attacks ($N=13$), where scammers use these advanced models to create legitimate-looking emails or SMS to earn the trust of potential victims. While GenAI models offer numerous benefits, these studies show that they pose significant risks when leveraged for malicious purposes, particularly in the realm of social engineering. GenAI can generate coherent, contextually relevant, and grammatically correct emails that mimic the style and tone of professional communication. This increases the likelihood of victims perceiving fraudulent emails as legitimate and trusting the message schmitt2023digital .
Regarding other scams, we found a variety of studies that try to detect Social Media Scams ($N=6$). These scams included a variety of fraudulent activities including fake user accounts and online groups, advertisement of fraudulent apps or fake phishing website aimed at stealing personal information or money.
@@
Similarly, $3$ studies focus on Romance Fraud via analysing profiles on social media and discussions with victims. We also identified $3$ studies that try to understand fraudulent cryptocurrency investment scams and $2$ studies that attempt to detect the likelihood of cryptocurrency manipulation. Finally, we identified $2$ studies that analysed fraudulent crowdfunding online and $1$ that studies fraudulent e-commerce websites.
@@ 4 Summary of Findings
We now provide a summary of our findings, categorized per fraud activity analyzed within the papers that are included in our SLR for qualitative analysis.
@@ 4.1 Data Sources
First, we report the most popular data sources used and the datasets analyzed.
@@
Phishing URLs. We start with understanding the preferable data sources and methodologies employed for the analysis and detection of Phishing URLs, as it is the most popular scam type category we detected in our SLR. In total, we analyze the data sources and detection methodologies of the identified $63$ papers that attempt to tackle this issue. Table LABEL:tbl:data_phishing_URL summarizes the data sources and methodologies used to detect Phishing URLs, as found in the literature.
@@
First, we find that researchers used various websites that offer information on URLs for the analysis of malicious and legitimate domains. This information may be webpage rankings (how trusted the webpage is), phishing reports, and historical data. By far, the most popular data source used was PhishTank, [^2] https://phishtank.org/ a website that allows users to report webpages that might be malevolent or suspicious, with $25$ studies using it as already labeled malicious websites li2022mui ; vo2023shark ; al2020convolutional ; aslam2023phish ; jishnu2023enhanced ; jaber2022improving ; rafsanjani2023qsecr ; alswailem2019detecting ; mandadi2022phishing ; jha2023intelligent ; marimuthu2022intelligent ; adebowale2023intelligent ; shaiba2022hunger ; rao2020two ; salloum2021phishing ; orunsolu2022predictive ; rao2019detection ; barraclough2021intelligent ; almseidin2019phishing ; chiew2019new ; li2019stacking ; sahingoz2019machine ; somesha2020efficient ; tharani2020understanding ; do2020malicious ; pradeepa2022lightweight .
@@
Another website that offers a list of phishing URLs is OpenPhish [^3] https://openphish.com/ and it was used by $3$ studies vo2023shark ; almseidin2019phishing ; chiew2019new . Two studies used URLhaus [^4] https://urlhaus.abuse.ch/, a project for sharing malicious URLs, for the collection and analysis of phishing URLs rafsanjani2023qsecr ; do2020malicious . We also find one work that used SpamHaus [^5] https://www.spamhaus.org/ for the collection of IP and domain reputation fernandez2022early , and one that used URLscan [^6] https://urlscan.io/ chen2021ai . Interestingly, chen2021ai also collected user-reported domains from ScammerInfo [^7] https://scammer.info/, a forum where users post and discuss various scams. Lastly, the webpage WhoIs [^8] https://who.is/, a webpage that offers historical data on webpages, was used for feature collection from shalke2022social ; adebowale2023intelligent .
@@
The most used data source for collection of legitimate webpages was Amazon Alexa, a webpage that offers webpage rankings, with $10$ studies using it to collect legitimate annotated webpages li2022mui ; navyah2022ensemble ; saha2023phishing ; liang2019using ; shaiba2022hunger ; orunsolu2022predictive ; rao2019detection ; somesha2020efficient ; tharani2020understanding ; do2020malicious ; liang2019using . Google’s Search Engine was used for one study chen2022development , while another used Majestic Million site [^9] https://majestic.com/reports/majestic-million for legitimate webpage collection jishnu2023enhanced , a site similar to Alexa.
@@
We also find that many records use previously publicly available datasets for their analysis. More specifically, $11$ studies gu2022ensemble ; jha2023machine ; jha2023machine ; mehndiratta2023malicious ; jain2023support ; saha2020phishing ; kumar2023hybrid ; ashwitha2023perception ; dr2023malicious ; pradeepa2022lightweight ; zamir2020phishing used already publicly available datasets published on Kaggle. Similarly, other $5$ studies navyah2022ensemble ; jaber2022improving ; kalabarige2023boosting ; mohammed2022accuracy ; priya2020gravitational used the UCI Phishing Dataset. [^10] https://archive.ics.uci.edu/dataset/327/phishing+websites Other studies used publicly available datasets from other sources ariyadasa2022phishrepo ; villanueva2022application ; vecile2022malicious ; kumar2020phishing ; zin2023machine ; pradeepa2022lightweight ; pathak2023classification .
@@
The most recent studies (published in 2023) that attempt to detect Phishing URLs automatically, attempt to do so by collecting data from alternative sources like social networks, and user-reported phishing URLs bitaab2023beyond ; saha2023phishing ; saha2023phishing ; bitaab2023beyond ; nakano2023canary ; janet2022real , while others used datasets from Telecom and Security institutions mehndiratta2023malicious ; yu2024efficient ; liang2019using Alas, we failed to detect the data source used for $9$ studies, as they do not clearly report how or from where they acquired the dataset they use for their study alkawaz2022identification ; el2021malweb ; kundra2023identification ; chen2020intelligent ; ou2021no ; raja2021lexical ; yadollahi2019adaptive ; nagy2023phishing ; geyik2021detection .
@@
Phishing Emails. We now discuss the data sources used in the $29$ works that tackle the Phishing Email detection problem. The data extracted from the literature and presented in this section are depicted in Table LABEL:tbl:data_phishing_emails.
@@
Two works genc2021understanding ; jiang2024detecting used emails received on the author’s personal or professional email spam folder. Other works that include datasets from alternative sources is the one of Mehdi et al. mehdi2023adversarial , which used various techniques to develop their dataset, like GPT2 generated synthetic phishing emails made available by previous research radford2019language , along with TextAttack [^11] https://github.com/QData/TextAttack, a Python framework for adversarial attacks, data augmentation, and model training in NLP, Textfooler [^12] https://github.com/jind11/TextFooler, a Model for Natural Language Attack on Text Classification, and PWWS ren2019generating . Another alternative data source for phishing email detection was used by Janez et al. janez2021trustworthiness who used data from SPAM Archive [^13] http://untroubled.org/spam/, a website that publishes spammy email repositories at the end of every month and is constantly updated.  gallo2019identifying used user reported e-mails. Lastly, the data source used by three works was not clearly stated within the manuscript mughaid2022intelligent ; venugopal2022detection ; rahmad2020performance .
@@
Phishing SMS. Regarding Phishing SMS (smishing), we included and analyzed $20$ papers in this SLR. For the detailed data, refer to Table LABEL:tbl:data_smishing.
@@
Similarly to previous analyses, we find that the overwhelming majority of works opted for using already publicly available datasets for their analysis and training a model to detect fraudulent SMS automatically. More specifically, a variety of subsets from a publicly available dataset on Kaggle [^14] https://www.kaggle.com/datasets/uciml/sms-spam-collection-dataset was used by $14$ studies vinothkumar2022detection ; agrawal2023effective ; jain2019feature ; mishra2023dsmishsms ; abid2022spam ; kohilan2023machine ; siagian2023improving ; gandhi2023sms ; al2023multi ; zhang2023bert ; dharani2023spam ; addanki2023safeguarding ; ulfath2021hybrid ; jain2022sms . Another study zhang2020lies used the Kaggle dataset, but incorporated Fake Base Station data and made it available to researchers. [^15] https://github.com/Cypher-Z/FBS_SMS_Dataset Similarly, this work vinothkumar2022detection used a subset of the Kaggle dataset in combination with emails and YouTube comments for spam content detection, while Lai et al. lai2022semorph used data provided by users. [^16] https://www.datafountain.cn/competitions/508 Tang et al. tang2022clues collected tweets where users were reporting smishing for their analysis. Two other works used data from the Korean Internet and Security Agency seo2024device and 360 Mobile Safe liu2021detecting . Lastly, Timko et al. timko2023commercial proposed a platform where users can freely post Phishing SMS for researchers to use. [^17] https://smishtank.com/
@@
Phishing Phone Calls. We identified $12$ studies that used phone call transcripts to understand Voice call-enabled phishing (vishing).
@@
Derakhshan et al. derakhshan2021detecting used the CallHome dataset which includes 120 unscripted 30-minute telephone conversations between native speakers of English. [^18] https://catalog.ldc.upenn.edu/LDC97S42 Another work djire2023evaluating used AI-generated deepfake voice recordings (Tacotron 2191919https://pytorch.org/hub/nvidia_deeplearningexamples_tacotron2/, Deepvoice 3202020https://r9y9.github.io/deepvoice3_pytorch/, and FastSpeech 2 ren2020fastspeech ). For authentic voice recordings they used the synplaflex dataset sini2018synpaflex which is a corpus of 87 hours of audiobooks in French.
@@
Various works used telecommunication operator datasets, like liang2023telecom using fraudulent caller IDs and phone transcripts huang2022learning from telecommunication operators in China, hu2023urf4cct used data from the Public Security Bureau in Zhejiang Province, China, and kim2021voice used data from the Korean Financial Supervisory Service. Kale et al. kale2021classification developed their dataset via questionnaires and victim testimonies. Other works collected data from various social networks, like YouTube transcripts rahman2022phone , Facebook, online blogs and forums, public datasets, as well as some that were developed based on studies of scammers’ activities and behaviors hong2023scam . Others opted for using previously analyzed and publicly available data gowri2021detection ; malhotra2023detection , while the data used from Zhong et al. zhong2020encoding was unclear.
User Reports. Four studies used user reports to understand phishing activities.
@@
First, liu2021fine constructed a fraud complaint dataset from the Internet finance service in China. Similarly, zhou2022keyword used court documents from Chinese online judgement records, while palad2019document used incident record forms from victims and interviews in the Philippines. Lastly, lwin2023supporting launch and introduce a website operated by the National Crime Prevention Council (NCPC) in Singapore, where users can report and get informed on the latest phishing activities.  [^21] https://www.scamalert.sg/
@@
Fake Reviews. We move to the Fake/Fraudulent Review scam type. We include $23$ studies in our analysis and the data extracted from each record in our analysis is depicted in Table LABEL:tbl:data_fake_reviews
@@
Other studies used application reviews from Google Play Store or Apple’s App Store yugeshwaran2022rank ; tushev2022domain ; obie2022violation . Other’s used previously available Amazon product reviews, [^24] https://snap.stanford.edu/data/web-Amazon.html or collected Amazon reviews rangar2022machine ; iqbal2023efficient ; furia2020tool ; gupta2020detecting ; thilagavathy2023fake ; chandana2021analyzing ; akshara2023small ; deekshan2022detection .
@@
Others collected Amazon Hotel and Holiday packages reviews rangari2022empirical ; silpa2023detection , or TripAdvisor review data tufail2022effect Lastly, bevendorff2024product used YouTube transcripts to interpret false review exaggeration. The source of the data used by Ganesh et al. ganesh2023implementation was unclear.
@@
Fraudulent Recruitment. We detected $19$ studies that focus on the detection of fraudulent job postings. The related data is listed in Table LABEL:tbl:data_recruitment.
@@
The overwhelming majority ($N=16$), of papers used the same publicly available dataset from Kaggle, [^25] https://www.kaggle.com/datasets/amruthjithrajvr/recruitment-scam, which holds about 18K job postings out of which 800 are fraudulent. Notably, this dataset includes data from 2012 to 2014 nessa2022recruitment ; prathaban2022verification ; pandey2022effective ; ranparia2020fake ; habiba2021comparative ; reddy2023web ; santhiya2023fake ; lal2019orfdetector ; nasser2021online ; amaar2022detection ; sofy2023intelligent ; vo2021dealing ; nanath2023investigation ; ullah2023smart ; bhatia2022detection ; li2021exploratory .
@@
The other three studies developed custom crawlers to collect data from various job posting sites in the UK (SEEK, Glassdoor, Indeed, and Gumtree) mahbub2022online , in Bangladesh (job.com.bd, bdjobstoday, deshijob) tabassum2021detecting , and in China (China–Boss, Zhipin, Liepin, and 51job) zhang2023orfpprediction .
@@
Fake Accounts Online. Various studies attempt to tackle the automated detection of fake profiles online. In Table LABEL:tbl:data_fake_users we depict the data extracted from the identified records.

### v2

@@ 3.1.3 Data extraction
Next, the authors agreed on the data to be extracted from the included studies. Only one of the three reviewers carried out the task of extracting data. This was deemed sufficient since the reviewer’s role involved extracting the required details from the papers, and a second reviewer did not have to check accuracy. The only aspect of the data extraction that the reviewer had to conceptualize was the specific Fraud Type analysed by the study under question. For example, if a study analysed Phishing URLs, it was labelled as Phishing URLs.
Not all papers explicitly specified the type of fraud analysed. Due to the diversity of scams, there is no agreed way of labelling fraud types. As a result, we used an umbrella term to categorise them. For example, online scam campaigns made by bot users on various social media platforms to advertise fraudulent phishing URLs include fake users and phishing URLs analysis; hence, we agreed to label papers of broad online scam campaigns as social media scams.
We did not classify a paper with more than one fraud type to ease the representation of our findings. Instead, the reviewer recorded the paper’s primary goal when identifying the fraud analysed. For example, if a paper used phishing emails to extract phishing URLs towards detecting phishing URLs, then that paper would be labelled as Phishing URLs, as that was the study’s main goal. The final version of the data extracted from each record is depicted in Table 2, along with relevant examples. A thematic analysis was conducted on the extracted data of the studies included for qualitative analysis, and we present our findings in Section 5.
@@ 4.1 Study selection and characteristics
The PRISMA-ScR flow diagram in Figure 3 summarises the study selection process. We identified $2,617$ studies for eligibility screening. The ACM Digital Library returned $389$ documents, IEEE Xplore returned $712$ documents, Web of Science returned $253$ documents, ProQuest returned $783$ documents, Google Scholar returned $399$ documents, and ArXiv returned $47$ documents. Experts in the area recommended an additional $34$ papers. After removing duplicates, $2,457$ papers remained for further review.
@@
At this stage, $10\%$ of these papers were selected ($N=242$) for Inter-Rater Reliability to calculate the multi-annotator agreement between the three annotators of this review. The Fleiss Kappa score between all three annotators was $0.83$, indicating almost perfect agreement. The Cohen Kappa Agreement was also calculated between each pair of annotators. The agreement between annotators AP and NT was $0.65$ (substantial agreement), between AP and EM was $0.66$, and between NT and EM was $0.52$ (moderate agreement). [^1] AP stands for author Antonis Papasavva, NT for author Nilufer Tuptuk, and EM for author Enrico Mariconti. The three annotators compared their annotation process and reviewed this SLR’s eligibility criteria and goals. Then, the lead annotator performed the rest of the annotations of the papers included in this review.
Reviewing the abstract of those papers resulted in $2,107$ papers being excluded from the study as they did not fit the eligibility criteria discussed in Section 3.1.2. This resulted in $350$ full-papers passing to eligibility screening, out of which $127$ did not fit the eligibility criteria and were excluded. Overall, this process resulted in $223$ full-text papers being included for qualitative analysis.
@@
4.2 Types of Online Fraud Identified in the Literature
@@
Figure 4 shows the types of fraud analysed in the full-text papers included in the qualitative analysis. The reviewer of these studies manually coded each paper with the scam type that the study focuses on, based on its title, abstract, and methodology. The majority of studies focus on phishing detection, with about a third ($29\%$) of the studies analysing phishing URLs online ($N=64$). More specifically, these works tackle the problem of automatically detecting whether a URL is likely fraudulent. Many papers were related to detecting phishing emails ($N=29$), followed by studies on SMS phishing detection ($N=20$). Other studies on phishing include phone call transcripts towards understanding and detecting voice call phishing ($N=12$), and a few studies attempt to understand phishing methods via victim reports ($N=4$).
@@
Moving on to other types of fraud, we found many studies that detect fake reviews on various platforms like the Google Play Store, Apple App Store, Yelp, and TripAdvisor ($N=23$). Another widely studied scam was recruitment fraud ($N=20$). We also found several studies that employed AI techniques to detect fake accounts on Facebook, Instagram, and Twitter ($N=18$).
@@
Similarly, $3$ studies focus on romance fraud via analysing profiles on social media and discussions with victims. We also identified $3$ studies that analysed and detected fraudulent cryptocurrency investment scams and $2$ studies that attempted to detect the likelihood of cryptocurrency manipulation. Finally, we identified $2$ studies that analysed fraudulent crowdfunding online and $1$ that studied fraudulent e-commerce websites.
Studies on other emerging types of fraud. A few studies have used Generative AI (GenAI) to analyse and better understand existing and emerging fraud techniques, especially ones where GenAI or LLMs are misused towards social engineering attacks. We identified $3$ studies that employed GenAI models to automatically interact with scammers to waste their resources and time while gathering information on the methods they used to defraud users, thereby disrupting their operations. This approach is defined as scam baiting: the process of using generative AI models to deceive and engage with scammers. Notably, this is a countermeasure against online fraud (“GenAI Scam Bait” in the Figure 4).
Our search also returned many studies that discuss the exploitation of GenAI towards social engineering attacks ($N=13$), where scammers use these advanced models to create legitimate-looking targeted emails or SMSs to earn the trust of potential victims. Although GenAI models offer numerous benefits, these studies show the significant risks they pose when used for malicious purposes, particularly in social engineering. GenAI can generate coherent, contextually relevant, and grammatically correct emails that mimic the style and tone of professional communication. This increases the likelihood of victims perceiving fraudulent emails as legitimate and trusting the message schmitt2023digital .
Regarding other scams, we found $6$ studies that detect social media scams. These scams included various fraudulent activities, like fake user accounts and online groups, advertisements of fraudulent apps, or phishing websites that aim to trick users into exposing their personal information or paying money.
The above three groups of studies were not identified in the literature discussed in Section 2; hence, we grouped and discussed them briefly here.
@@ 5 Summary of Findings
This section summarises our findings, categorized per fraud activity analysed within the papers included in our SLR for qualitative analysis.
@@ 5.1 Data Sources
First, we report the most popular data sources used, and the datasets analysed.
@@
Data Used for Phishing URL Detection. We start by understanding the chosen data sources for analysing and detecting phishing URLs; the most popular scam-type category we have detected in our SLR. We analysed the data sources and detection methodologies of the identified $63$ papers that focused on this issue. Table LABEL:tbl:data_phishing_URL summarizes these.
@@
Researchers used various websites that offer information on URLs for the analysis of malicious and legitimate domains. This information may be web page rankings (how trusted the webpage is), phishing reports, and historical data. By far, the most popular data source used was PhishTank [^2] https://phishtank.org/, a website that allows users to report webpages that might be malevolent or suspicious, with $25$ studies using it as already labelled malicious websites li2022mui ; vo2023shark ; al2020convolutional ; aslam2023phish ; jishnu2023enhanced ; jaber2022improving ; rafsanjani2023qsecr ; alswailem2019detecting ; mandadi2022phishing ; jha2023intelligent ; marimuthu2022intelligent ; adebowale2023intelligent ; shaiba2022hunger ; rao2020two ; salloum2021phishing ; orunsolu2022predictive ; rao2019detection ; barraclough2021intelligent ; almseidin2019phishing ; chiew2019new ; li2019stacking ; sahingoz2019machine ; somesha2020efficient ; tharani2020understanding ; do2020malicious ; pradeepa2022lightweight .
@@
Another website that offers a list of phishing URLs is OpenPhish [^3] https://openphish.com/ and it was used in $3$ studies vo2023shark ; almseidin2019phishing ; chiew2019new . Two studies used URLhaus [^4] https://urlhaus.abuse.ch/, a project for sharing malicious URLs, for the collection and analysis of phishing URLs rafsanjani2023qsecr ; do2020malicious . We also find one study that used SpamHaus [^5] https://www.spamhaus.org/ for the collection of IP and domain reputation fernandez2022early , and one that used URLscan [^6] https://urlscan.io/ chen2021ai . Interestingly, a study  chen2021ai also collected user-reported domains from ScammerInfo [^7] https://scammer.info/, a forum where users post and discuss various scams. Lastly, the webpage WhoIs [^8] https://who.is/, a webpage that offers historical data on webpages, was used for feature collection in two studies shalke2022social ; adebowale2023intelligent .
@@
The most used data source for the collection of legitimate webpages was “Alexa,” (a global ranking system that estimated a website’s popularity that shut down in May 2022) with $10$ studies using it to collect legitimate annotated webpages li2022mui ; navyah2022ensemble ; saha2023phishing ; liang2019using ; shaiba2022hunger ; orunsolu2022predictive ; rao2019detection ; somesha2020efficient ; tharani2020understanding ; do2020malicious . Google’s search engine was used for one study chen2022development , while another used the Majestic Million site [^9] https://majestic.com/reports/majestic-million for legitimate webpage collection jishnu2023enhanced , a site similar to Alexa.
@@
Many studies used existing publicly available datasets for their analysis. More specifically, $11$ studies gu2022ensemble ; jha2023machine ; jha2023machine ; mehndiratta2023malicious ; jain2023support ; saha2020phishing ; kumar2023hybrid ; ashwitha2023perception ; dr2023malicious ; pradeepa2022lightweight ; zamir2020phishing used publicly available datasets published on Kaggle (a repository for researchers to publish data). Similarly, $5$ studies navyah2022ensemble ; jaber2022improving ; kalabarige2023boosting ; mohammed2022accuracy ; priya2020gravitational used the UCI Phishing Dataset. [^10] https://archive.ics.uci.edu/dataset/327/phishing+websites Other studies used publicly available datasets from other sources ariyadasa2022phishrepo ; villanueva2022application ; vecile2022malicious ; kumar2020phishing ; zin2023machine ; pradeepa2022lightweight ; pathak2023classification .
@@
The most recent studies (published in 2023) that attempted to detect phishing URLs automatically collected data from alternative sources like social networks and user-reported phishing URLs bitaab2023beyond ; saha2023phishing ; saha2023phishing ; bitaab2023beyond ; nakano2023canary ; janet2022real , while others used datasets from telecom and Security organizations mehndiratta2023malicious ; yu2024efficient ; liang2019using . Alas, we failed to detect the data source used by $9$ studies, as the authors did not report how or from where they acquired the dataset used in their study alkawaz2022identification ; el2021malweb ; kundra2023identification ; chen2020intelligent ; ou2021no ; raja2021lexical ; yadollahi2019adaptive ; nagy2023phishing ; geyik2021detection .
@@
Data Used for Phishing Email Detection. We now discuss the data sources used in the $29$ works that tackle phishing email detection. The data extracted from the literature and presented in this section are shown in Table LABEL:tbl:data_phishing_emails.
@@
Two studies genc2021understanding ; jiang2024detecting used emails received in the author’s personal or professional email spam folder. Another study that included datasets from alternative sources was by Mehdi et al. mehdi2023adversarial . They used various techniques to develop their dataset, including GPT2 generated synthetic phishing emails made available in previous research radford2019language , along with TextAttack [^11] https://github.com/QData/TextAttack, a Python framework for adversarial attacks, data augmentation, and model training in NLP, Textfooler [^12] https://github.com/jind11/TextFooler, a Model for Natural Language Attack on Text Classification, and Probability Weighted Word Saliency (PWWS) ren2019generating , a technique for generating adversarial text.
Another alternative data source for phishing email detection was used by Janez et al. janez2021trustworthiness who used data from SPAM Archive [^13] http://untroubled.org/spam/, a website that publishes spam email repositories at the end of every month and is constantly updated. Gallo et al. gallo2019identifying analysed user-reported emails. Lastly, the data source used in $3$ studies was not clearly stated within the manuscript mughaid2022intelligent ; venugopal2022detection ; rahmad2020performance .
@@
Data Used for Phishing SMS Detection. Regarding Phishing SMS (smishing), we included and analysed $20$ papers in this SLR. For the detailed data, refer to Table LABEL:tbl:data_smishing.
@@
Similarly to previous analyses, the overwhelming majority of works opted for using already publicly available datasets to analyse and train a model. More specifically, a variety of subsets from a publicly available dataset on Kaggle [^14] https://www.kaggle.com/datasets/uciml/sms-spam-collection-dataset was used by $14$ studies vinothkumar2022detection ; agrawal2023effective ; jain2019feature ; mishra2023dsmishsms ; abid2022spam ; kohilan2023machine ; siagian2023improving ; gandhi2023sms ; al2023multi ; zhang2023bert ; dharani2023spam ; addanki2023safeguarding ; ulfath2021hybrid ; jain2022sms . Another study zhang2020lies used the Kaggle dataset but incorporated Fake Base Station data and made it available to researchers. [^15] https://github.com/Cypher-Z/FBS_SMS_Dataset Similarly, this work vinothkumar2022detection used a subset of the Kaggle dataset in combination with emails and YouTube comments for spam content detection, while Lai et al. lai2022semorph used data provided by users. [^16] https://www.datafountain.cn/competitions/508 Tang et al. tang2022clues collected tweets where users reported smishing for their analysis. Two other works used data from the Korean Internet and Security Agency seo2024device and 360 Mobile Safe liu2021detecting . Lastly, Timko et al. timko2023commercial proposed a platform where users can freely post Phishing SMS for researchers to use. [^17] https://smishtank.com/
@@
Data Used for Phishing Phone Call Detection. We identified $12$ studies that used phone call transcripts to understand voice call-enabled phishing (vishing).
@@
Derakhshan et al. derakhshan2021detecting used the CallHome dataset, which includes 120 unscripted 30-minute telephone conversations between native speakers of English. [^18] https://catalog.ldc.upenn.edu/LDC97S42 Another study djire2023evaluating used AI-generated deepfake voice recordings (Tacotron 2191919https://pytorch.org/hub/nvidia_deeplearningexamples_tacotron2/, Deepvoice 3202020https://r9y9.github.io/deepvoice3_pytorch/, and FastSpeech 2 ren2020fastspeech ). For authentic voice recordings, they used the synplaflex dataset sini2018synpaflex , a corpus of audiobooks in French.
@@
Some studies used telecommunication operator datasets, like liang2023telecom using fraudulent caller IDs and phone transcripts huang2022learning from telecommunication operators in China, Hu and Yuan hu2023urf4cct used data from the Public Security Bureau in Zhejiang Province, China, and kim2021voice used data from the Korean Financial Supervisory Service. Kale et al. kale2021classification developed their dataset via questionnaires and victim testimonies. Other authors collected data from various social networks, including YouTube transcripts rahman2022phone , Facebook, online blogs and forums, public datasets, as well as some that were developed based on studies of scammers’ activities and behaviours hong2023scam . Others opted for using previously analysed and publicly available data gowri2021detection ; malhotra2023detection , while the data Zhong et al. zhong2020encoding used was unclear.
@@
Data Used for Phishing (User Reports) Detection. Four studies used user reports to understand phishing activities. First, one study liu2021fine constructed a fraud complaint dataset from the Internet finance service in China. Similarly, another zhou2022keyword used court documents from Chinese online judgement records, while a third palad2019document used incident record forms from victims and interviews in the Philippines. In the final study, the authors lwin2023supporting launched and introduced a website operated by the National Crime Prevention Council (NCPC) in Singapore, where users could report and receive information on the latest phishing activities.  [^21] https://www.scamalert.sg/
@@
Data Used for Fake Review Detection. For this type of scam, $23$ studies were included in our analysis, and the data extracted from them is depicted in Table LABEL:tbl:data_fake_reviews
@@
Other studies used application reviews from the Google Play Store or Apple’s App Store yugeshwaran2022rank ; tushev2022domain ; obie2022violation . Other studies used previously available Amazon product reviews [^24] https://snap.stanford.edu/data/web-Amazon.html, or collected Amazon reviews rangar2022machine ; iqbal2023efficient ; furia2020tool ; gupta2020detecting ; thilagavathy2023fake ; chandana2021analyzing ; akshara2023small ; deekshan2022detection .
@@
Some studies collected reviews of Amazon Hotel and Holiday packages  rangari2022empirical ; silpa2023detection , or TripAdvisor review data tufail2022effect And, lastly, one study bevendorff2024product used YouTube transcripts to interpret false review exaggeration. The data source used by Ganesh et al. ganesh2023implementation was unclear.
@@
Data Used for Recruitment Fraud Detection. We found $19$ studies that focused on the detection of fraudulent job postings (see Table LABEL:tbl:data_recruitment).
@@
The overwhelming majority ($N=16$) of papers, used the same publicly available dataset from Kaggle, [^25] https://www.kaggle.com/datasets/amruthjithrajvr/recruitment-scam, which holds about 18K job postings of which 800 are fraudulent. Notably, this dataset includes data from 2012 to 2014 nessa2022recruitment ; prathaban2022verification ; pandey2022effective ; ranparia2020fake ; habiba2021comparative ; reddy2023web ; santhiya2023fake ; lal2019orfdetector ; nasser2021online ; amaar2022detection ; sofy2023intelligent ; vo2021dealing ; nanath2023investigation ; ullah2023smart ; bhatia2022detection ; li2021exploratory .
@@
The other three studies developed custom crawlers to collect data from various job posting sites in the UK (SEEK, Glassdoor, Indeed, and Gumtree) mahbub2022online , in Bangladesh (job.com.bd, bdjobstoday, deshijob) tabassum2021detecting , and in China (China-Boss, Zhipin, Liepin, and 51job) zhang2023orfpprediction .
@@
Data Used for Fake Account Detection. Some studies attempted to tackle the automated detection of fake profiles online, and Table LABEL:tbl:data_fake_users shows the data extracted from them.